# Regularized Q-Learning with Linear Function Approximation

### v1

@@ 5 Finite-Time Guarantees
In this section, we present the finite-time convergence guarantee for Algorithm 1 and the performance bound of the derived policies.
@@ 5.1 Convergence Analysis
Many current gradient-based algorithms depend on the objective function’s smooth property, a characteristic not present in problem (3). We address this by considering a relaxed version of smoothness.
Lemma 5.1
(Relaxation of smoothness.) Suppose Assumptions 3.1, 3.2 hold, for all $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, then
@@
We refer the readers to Appendix 10.1 for detailed proof. A similar property termed $(L_{0},L_{1})$-smoothness is incorporated into the optimization algorithm, enhancing the speed of resolution, as demonstrated in [57, 58]. By employing the smooth truncation operator to impose bounds and integrating the projection operation as illustrated in equation (11a), we ensure that the $\ell_{2}$-norm of the stochastic gradient of the lower-level problem is constrained.
Lemma 5.2
@@
The detailed proof can be found in Appendix 10.2.
@@
Theorem 5.3
Suppose Assumptions 3.1, 3.2 and 3.4 hold. Let $\alpha=\alpha_{0}T^{-3/4},\beta=\beta_{0}T^{-1/2}$ and $\alpha<\frac{\sigma_{\min}^{2}}{2(L_{1}+\sigma_{\min})}$, where $\alpha_{0},\beta_{0}$ are positive constants. Then, Algorithm 1 yields
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|=\mathcal{O}\left(T^{-1/4}\right)$ (12a)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|=\mathcal{O}\left(T^{-1/4}\right).$ (12b)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|=\mathcal{O}\left(T^{-1/4}\right).$ (13)
The detailed proof can be found in Appendix 9.1.
Observe that the bounds outlined in Theorem 5.3 are expressed in terms of $\ell_{2}$ norms. This is different from the conventional use of squared $\ell_{2}$ norms prevalent in the majority of two-timescale convergent algorithms, as demonstrated in [55, 52, 31, 28, 59, 29, 30]. The primary reason for this distinction lies in the normalization of the gradient employed in update (11b), which is common in the optimization techniques like normalized or clipped gradient-based algorithms. However, it can be inferred that the norm scaling at an order of $T^{-1/4}$ is roughly equivalent to the squared norm scaling at an order of $T^{-1/2}$, albeit with a slight discrepancy.
Remark 5.4
The rates in (12a) and (13) have components that are inversely proportional to the square of $\lambda_{g}$, the lower bound of the eigenvalue of $\Sigma$. This suggests that a larger $\lambda_{g}$ leads to faster convergence of the lower-level problem, thereby enhancing the convergence of the upper-level problem. Additionally, a high value of $\lambda_{g}$, along with a large coefficient $\tau$, augments the smoothness of the objective function, as established in Lemma 5.1. Furthermore, the rates in (12b) and (13) include components inversely proportional to $\sigma_{\min}$, the lower bound of the singular value of the matrix $\Sigma_{\tau,\delta,\theta}$. Increased smoothness in $J$, together with a large $\sigma_{\min}$, allows for a greater step size $\alpha$. This, in practice, accelerates convergence. The effect of the threshold $\delta$ on the rates in Theorem 5.3 is more complex, and we defer the discussion to Section 6.
Remark 5.5
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\theta^{t}-\omega^{t}\right\|=\mathcal{O}\left(T^{-1/4}\right).$
The detailed proof is in Appendix 11.1.
@@ 5.2 Performance Analysis
In this subsection, we analyze the difference between the estimated state-action value function $\hat{Q}_{\theta}$ and the optimal value function $Q^{*}_{\tau}$ in $\mathcal{M}_{\tau}$. Additionally, we compare the performance of policies derived from Algorithm 1 against the optimal policies $\pi^{*}_{\tau}$.
@@
The performance of a policy $\pi$ is measured by $\mathbb{E}_{s\sim\mu_{0}}\left[V_{\pi,\tau}(s)\right]$ in regularized MDP $\mathcal{M}_{\tau}$, recall the definition of $V_{\pi,\tau}$ in (2.1). This expression quantifies the expected cumulative rewards and regularization terms starting from the initial state. Consequently, $\mathbb{E}_{s\sim\mu_{0}}\left[V_{\tau}^{*}(s)-V_{\pi_{\theta^{t}},\tau}(s)\right]$ captures the expected improvement achievable when transitioning from policy $\pi_{\theta^{t}}$ to the optimal policy $\pi_{\tau}^{*}$. To quantify this disparity in performance, we employ the $\ell_{\infty}$-norm which offers uniform upper bounds on the expected improvements. In the following theorem, we present the finite-time bound characterizing the difference between the estimated value function and the optimal value functions, and the performance guarantee for the sequence of policies $\left\{\pi_{\theta^{t}}\right\}_{t=0}^{T-1}$ produced by Algorithm 1.
Theorem 5.7
Suppose Assumptions 3.1, 3.2 and 3.4 hold. Then Algorithm 1 yields
@@
$\leq$ $\frac{1}{\sigma_{\min}T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|+\mathcal{E}_{\mathrm{approx}}+\gamma\left(\delta_{0}-\mathcal{K}_{\delta}\left(\delta_{0}\right)\right),$ (15)
and
$\frac{1-\gamma}{2T}$ $\sum_{t=0}^{T-1}\mathbb{E}\left\|V_{\tau}^{*}-V_{\pi_{\theta^{t}},\tau}\right\|_{\infty}$
$\leq$ $\frac{1}{\sigma_{\min}T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|+\mathcal{E}_{\mathrm{approx}}+\gamma\left(\delta_{0}-\mathcal{K}_{\delta}\left(\delta_{0}\right)\right),$ (16)
The detailed proof can be found in Appendix 9.2.
@@
The second bound in Theorem 5.7 establishes, to the best of our knowledge, the first performance guarantee of the derived policies for a convergent algorithm that learns the solution of the regularized MDP with linear function approximation. As the approximation error nears zero and the effect of the smooth truncation diminishes, the learned policy $\pi_{\theta^{t}}$ converges towards the optimal policy $\pi^{*}_{\tau}$. This convergence is realized under the condition of Bellman Completeness [61], provided that the threshold $\delta$ is also sufficiently large.
@@ 7.2 MountainCar-v0
We evaluated the policies derived from Algorithm 1 with a negative entropy regularizer and compared them with those obtained through Q-learning [14], Coupled Q-learning (CQL) [25], Greedy-GQ [27], and Double Q-learning (Double QL) [35]. We selected the coefficient of the regularizer for Algorithm 1 from the set $\{0.01,0.05\}$. For Algorithm 1, we adopted the regularized policy $\pi_{\theta}$ defined in 4 as the behavioral policy. In contrast, an $\epsilon$-greedy policy, with $\epsilon=0.1$, served as the behavioral policy for the other mentioned algorithms.
@@
[…]
8 Auxiliary Lemmas
Lemma H1
Under Assumptions 3.1 and 3.2, $\omega^{*}(\cdot)$ is Lipschitz with respect to $\theta\in\mathbb{R}^{d}$ with constant $\gamma/\lambda_{g}$.
Proof 8.2.
The gradient of $\omega^{*}(\theta)$ with respect to $\theta\in\mathbb{R}^{d}$ is {IEEEeqnarray*}rl &_θω^*(θ)
= γΣ^-1
E_D[z(s^;τ, δ,θ)_a^′∈Aπ_θ(a^′∣s^)ϕ(s,a) ϕ(s^,a^)^⊤]. The $\ell_{2}$-norm of the gradient of $\omega^{*}(\theta)$ is bounded as $\left\|\nabla_{\theta}\omega^{*}(\theta)\right\|\leq 1/\lambda_{g}$. According to the mean value theorem, we have {IEEEeqnarray*}rCr ∥ω^*(θ_1)-ω^*(θ_2)∥ &≤max_θ ∥_θω^*(θ)∥∥θ_1-θ_2∥
1λg ∥θ_1-θ_2∥.
Lemma 8.3.
@@ 8 Auxiliary Lemmas
Proof 8.4.
By the properties of the regularizer and Assumption 3.2, for $s\in\mathcal{S}$ and $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, we have {IEEEeqnarray*}rlr ∥π_θ_1(s)- π_θ_2( s)∥ =& ∥G^*_τ(^Q_θ_1(s,⋅))-G^*_τ(^Q_θ_2(s,⋅)) ∥
LGτ ∥^Q_θ_1(s,)-^Q_θ_2(s,)∥
LGA—τ ∥θ_1-θ_2∥.
Lemma 8.5.
@@
$\nabla_{\theta}\hat{G}^{2}_{\tau,\delta}$ $(\hat{Q}_{\theta}(s^{\prime},\cdot))$
$=$ $\nabla_{\theta}\mathcal{K}^{2}_{\delta}(G^{*}_{\tau}(\hat{Q}_{\theta}(s^{\prime},\cdot)))$
@@
[…]
$\leq\left\|\mathbb{E}_{\mathcal{D}}\left[\phi\left(s^{\prime},\cdot\right)\left(\pi_{\theta_{1}}\left(\cdot\mid s^{\prime}\right)-\pi_{\theta_{2}}\left(\cdot\mid s^{\prime}\right)\right)\phi(s,a)^{\top}\right]\right\|$
@@
where the last inequality is due to Lemma 8.3 and (18).
Lemma 8.7.
@@
$\left\|\theta^{t+1}-\theta^{t}\right\|$
$\leq\alpha\left\|\gamma z\left(s^{\prime};\tau,\delta,\theta^{t}\right)\sum_{a^{\prime}\in\mathcal{A}}\!\!\pi_{\theta^{t}}\left(a^{\prime}\mid s^{\prime}_{t}\right)\phi(s_{t}^{\prime},a^{\prime})-\phi(s_{t},a_{t})\right\|$
$\leq 2\alpha.$
Lemma 8.9.
@@
Proof 8.10.
Given the definitions of the gradient $\nabla J(\theta)$ and the surrogate gradient ¯$\bar{\nabla}_{\theta}f(\theta,\omega)$ in (3.2) and (10b), respectively, we have
@@
9 Proofs of Theorems
9.1 Proof of Theorem 5.3
Let $\mathcal{F}_{t}:=\sigma\left\{\omega^{0},\theta^{0},\ldots,\omega^{t},\theta^{t}\right\}$ and $\mathcal{F}^{\prime}_{t}:=\sigma\left\{\omega^{0},\theta^{0},\ldots,\omega^{t},\theta^{t},\omega^{t+1}\right\}$ be the filtration of the random variables up to iteration $t$, where $\sigma\{\cdot\}$ denotes the $\sigma$-algebra generated by the random variables. Similar to the definition in (10a) and (10b), suppose $(\tilde{s},\tilde{a},\tilde{s}^{\prime})\sim\mathcal{D}$, we define $\tilde{h}_{g}^{t}$ and $\tilde{h}_{f}^{t}$ as:
$\tilde{h}_{g}^{t}$ $\triangleq\phi(\tilde{s},\tilde{a})\Big{(}\phi(\tilde{s},\tilde{a})^{\top}\omega^{t}-R(\tilde{s},\tilde{a})-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta^{t}}\left(\tilde{s}^{\prime}_{t},\cdot\right))\Big{)},$
$\tilde{h}_{f}^{t}$ $\triangleq\bigg{(}\gamma z\left(\tilde{s}^{\prime};\tau,\delta,\theta^{t}\right)\sum_{\tilde{a}^{\prime}\in\mathcal{A}}\pi_{\theta^{t}}\left(\tilde{a}^{\prime}\mid\tilde{s}^{\prime}\right)\phi(\tilde{s}^{\prime},\tilde{a}^{\prime})-\phi(\tilde{s},\tilde{a})\bigg{)}$
×$\qquad\qquad\qquad\ \ \ \ \ \ \ \ \ \ \ \ \times\phi(\tilde{s},\tilde{a})^{\top}\left(\omega^{t+1}-\theta^{t}\right).$
It follows that $\mathbb{E}[\tilde{h}_{g}^{t}\mid\mathcal{F}_{t}]=\nabla_{\omega}g(\theta^{t},\omega^{t})$ and $\mathbb{E}[\tilde{h}_{f}^{t}\mid\mathcal{F}_{t}^{\prime}]=\bar{\nabla}_{\omega}f\left(\theta^{t},\omega^{t+1}\right)$. Also, under Assumption 3.1, we define $T_{\alpha}$ as the number of iterations needed to ensure the bias incurred by the Markovian sampling is sufficiently small:

### v3

@@ 5 Finite-Time Guarantees
In this section, we present the finite-time convergence guarantee for Algorithm 1 and a finite-time performance bound of the derived policies. The approach we take is similar to that of [41, 14] wherein the two time scales refer to different constant step sizes for lower level and upper level updates, i.e. $\alpha<\beta$ and the finite-time guarantee is expressed in terms of a target number of iterations $T\geq 1$. This differs from other approaches in which the step-sizes converge to zero at different rates as in [27, 42, 43].
@@ 5.1 Convergence Analysis
The upper level objective function in (5) is not smooth. We consider the following relaxation of smoothness:
Lemma 5.1.
(Relaxation of smoothness.) Suppose Assumptions 2.2, 2.3 hold, for all $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, then
@@
We refer the readers to Appendix C.2 for detailed proof. A similar property termed $(L_{0},L_{1})$-smoothness is incorporated into the optimization algorithm, enhancing the speed of resolution, as demonstrated in [44, 45]. By employing the smooth truncation operator to impose bounds and integrating the projection operation as illustrated in equation (12a), we ensure that the $\ell_{2}$-norm of the stochastic gradient of the lower-level problem is constrained.
Lemma 5.2.
@@
The detailed proof can be found in Appendix C.3.
@@
Theorem 5.3.
Suppose Assumptions 2.2, 2.3 and 3.2 hold. Let $\alpha=\alpha_{0}T^{-3/4},\beta=\beta_{0}T^{-1/2}$, where $\alpha_{0},\beta_{0}$ are positive constants, and
$\alpha\leq\frac{\beta}{\frac{\log(\beta/\kappa)}{\log\rho}+2},\beta\leq$ $\frac{\sigma_{\min}}{4},$
$64\beta+16\beta^{2}+8\alpha^{2}\leq$ $\frac{\sigma_{\min}^{2}}{L_{1}}.$
Then, Algorithm 1 yields
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right)$ (13a)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$ (13b)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$ (14)
The detailed proof can be found in Appendix B.1.
Observe that the bounds outlined in Theorem 5.3 are expressed in terms of $\ell_{2}$ norms. This is different from the conventional use of squared $\ell_{2}$ norms prevalent in the majority of two-timescale convergent algorithms, as demonstrated in [14, 39, 33, 25, 46, 32, 31]. The primary reason for this distinction lies in the normalization of the gradient employed in update (12b), which is common in the optimization techniques like normalized or clipped gradient-based algorithms. However, it can be inferred that the norm scaling at an order of $T^{-1/4}(\log T)^{1/2}$ is roughly equivalent to the squared norm scaling at an order of $T^{-1/2}\log T$, albeit with a slight discrepancy.
Remark 5.4.
The rates in (13a) and (14) have components that are inversely proportional to the square of $\lambda_{g}$, the lower bound of the eigenvalue of $\Sigma$. This suggests that a larger $\lambda_{g}$ leads to faster convergence of the lower-level problem, thereby enhancing the convergence of the upper-level problem. Additionally, a high value of $\lambda_{g}$, along with a large coefficient $\tau$, augments the smoothness of the objective function, as established in Lemma 5.1. Furthermore, the rates in (13b) and (14) include components inversely proportional to $\sigma_{\min}$, the lower bound of the singular value of the matrix $\hat{\Sigma}_{\tau,\delta,\theta}$. Increased smoothness in $J$, together with a large $\sigma_{\min}$, allows for a greater step size $\alpha$. This, in practice, accelerates convergence. The effect of the threshold $\delta$ on the rates in Theorem 5.3 is more complex, and we defer the discussion to Section 6.
Remark 5.5.
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\theta^{t}-\omega^{t}\right\|=\mathcal{O}\left(T^{-1/4}(\log T)^{1/2}\right).$
The detailed proof is in Appendix D.1.
@@ 5.2 Performance Analysis
In this subsection, we analyze the difference between the estimated state-action value function $\hat{Q}_{\theta}$ and the optimal value function $Q^{*}_{\tau}$ in $\mathcal{M}_{\tau}$.
@@
In the following theorem, we present the finite-time bound characterizing the difference between the estimated value function and the optimal value functions.
Theorem 5.7.
Suppose Assumptions 2.2, 2.3 and 3.2 hold. Then Algorithm 1 yields
@@
$\leq$ $\frac{1}{\sigma_{\min}T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|+\mathcal{E}_{\mathrm{approx}}+\gamma\left(\delta_{0}-\mathcal{K}_{\delta}\left(\delta_{0}\right)\right).$ (16)
The detailed proof can be found in Appendix B.2.
@@
As the approximation error nears zero and the effect of the smooth truncation diminishes, the learned policy $\pi_{\theta^{t}}$ converges towards the optimal policy $\pi^{*}_{\tau}$. This convergence is realized under the condition of Bellman Completeness [48], provided that the threshold $\delta$ is also sufficiently large.
A corollary to Theorem 5.7 characterizes policy convergence by a straightforward application of Proposition 2.1(iii):
Corollary 5.8.
Suppose Assumptions 2.2, 2.3 and 3.2 hold. Then:
$\frac{1-\gamma}{T}\!\sum_{t=0}^{T-1}\!\mathbb{E}\left\|\pi^{*}_{\tau}-\pi_{\theta^{t}}\right\|_{\infty}$ $\!\leq\!\frac{\sqrt{|\mathcal{A}|}L_{G}}{\tau}\Big{(}\!\frac{1}{\sigma_{\min}T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla J\left(\theta^{t}\right)\right\|$
$+\mathcal{E}_{\mathrm{approx}}+\gamma\left(\delta_{0}-\mathcal{K}_{\delta}\left(\delta_{0}\right)\right)\Big{)}.$
The detailed proof is in Appendix D.2.
@@ 7.2 MountainCar-v0
We evaluated the policies derived from Algorithm 1 with a negative entropy regularizer and compared them with those obtained through Q-learning [50], Coupled Q-learning (CQL) [26], Greedy-GQ [20], and Double Q-learning (Double QL) [18]. We selected the coefficient of the regularizer for Algorithm 1 from the set $\{0.01,0.05\}$. For Algorithm 1, we adopted the regularized policy $\pi_{\theta}$ defined in 3 as the behavioral policy. In contrast, an $\epsilon$-greedy policy, with $\epsilon=0.1$, served as the behavioral policy for the other mentioned algorithms.
@@
Appendix A Auxiliary Lemmas
Lemma A1.
Under Assumptions 2.2 and 2.3, $\omega^{*}(\cdot)$ is Lipschitz with respect to $\theta\in\mathbb{R}^{d}$ with constant $\gamma/\lambda_{g}$.
Proof.
The gradient of $\omega^{*}(\theta)$ with respect to $\theta\in\mathbb{R}^{d}$ is {IEEEeqnarray*}rl ∇_θω^*(θ)=γΣ^-1E_D[z(s^′;τ, δ,θ)∑_a^′∈Aπ_θ(a^′∣s^′)ϕ(s,a) ϕ(s^′,a^′)^⊤]. The $\ell_{2}$-norm of the gradient of $\omega^{*}(\theta)$ is bounded as $\left\|\nabla_{\theta}\omega^{*}(\theta)\right\|\leq 1/\lambda_{g}$. According to the mean value theorem, we have {IEEEeqnarray*}rCr ∥ω^*(θ_1)-ω^*(θ_2)∥ ≤max_θ ∥∇_θω^*(θ)∥∥θ_1-θ_2∥≤1λgθ_1-θ_2∥. ∎
Lemma A2.
@@ Appendix A Auxiliary Lemmas
Proof.
By the properties of the regularizer and Assumption 2.3, for $s\in\mathcal{S}$ and $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, we have {IEEEeqnarray*}rlr ∥π_θ_1(⋅∣s)- π_θ_2(⋅∣ s)∥ =& ∥∇G^*_τ(^Q_θ_1(s,⋅))-∇G^*_τ(^Q_θ_2(s,⋅)) ∥
≤ LGτ ∥^Q_θ_1(s,⋅)-^Q_θ_2(s,⋅)∥
≤ LG|A|τθ_1-θ_2∥. ∎
Lemma A3.
@@
$\nabla_{\theta}\hat{G}^{2}_{\tau,\delta}(\hat{Q}_{\theta}(s^{\prime},\cdot))$ $=\nabla_{\theta}\mathcal{K}^{2}_{\delta}(G^{*}_{\tau}(\hat{Q}_{\theta}(s^{\prime},\cdot)))$
@@
$\|\hat{\Sigma}_{\tau,\delta,\theta_{1}}-\hat{\Sigma}_{\tau,\delta,\theta_{2}}\|$ $\leq\left\|\mathbb{E}_{\mathcal{D}}\left[\phi\left(s^{\prime},\cdot\right)\left(\pi_{\theta_{1}}\left(\cdot\mid s^{\prime}\right)-\pi_{\theta_{2}}\left(\cdot\mid s^{\prime}\right)\right)\phi(s,a)^{\top}\right]\right\|$
@@
where the last inequality is due to Lemma A2 and (18). ∎
Lemma A4.
@@
$\left\|\theta^{t+1}-\theta^{t}\right\|\leq\alpha\left\|\gamma z\left(s^{\prime};\tau,\delta,\theta^{t}\right)\sum_{a^{\prime}\in\mathcal{A}}\!\!\pi_{\theta^{t}}\left(a^{\prime}\mid s^{\prime}_{t}\right)\phi(s_{t}^{\prime},a^{\prime})-\phi(s_{t},a_{t})\right\|\leq 2\alpha.$
Lemma A5.
@@
Proof.
Given the definitions of the gradient $\nabla J(\theta)$ and the surrogate gradient ¯$\bar{\nabla}_{\theta}f(\theta,\omega)$ in (3.3) and (11), respectively, we have
@@
Appendix B Proofs of Theorems
B.1 Proof of Theorem 5.3
Let $\mathcal{F}_{t}:=\sigma\left\{\omega^{0},\theta^{0},\ldots,\omega^{t},\theta^{t}\right\}$ and $\mathcal{F}^{\prime}_{t}:=\sigma\left\{\omega^{0},\theta^{0},\ldots,\omega^{t},\theta^{t},\omega^{t+1}\right\}$ be the filtration of the random variables up to iteration $t$, where $\sigma\{\cdot\}$ denotes the $\sigma$-algebra generated by the random variables. Under Assumption 2.2, we define $T_{\alpha}$ as the number of iterations needed to ensure the bias incurred by the Markovian sampling is sufficiently small: